### v2

# On the Convergence of Zeroth-Order Federated Tuning for Large Language Models

Abstract.
The confluence of Federated Learning (FL) and Large Language Models (LLMs) is ushering in a new era in privacy-preserving natural language processing. However, the intensive memory requirements for fine-tuning LLMs pose significant challenges, especially when deploying on clients with limited computational resources. To circumvent this, we explore the novel integration of Memory-efficient Zeroth-Order Optimization within a federated setting, a synergy we term as FedMeZO. Our study is the first to examine the theoretical underpinnings of FedMeZO in the context of LLMs, tackling key questions regarding the influence of large parameter spaces on optimization behavior, the establishment of convergence properties, and the identification of critical parameters for convergence to inform personalized federated strategies. Our extensive empirical evidence supports the theory, showing that FedMeZO not only converges faster than traditional first-order methods such as FedAvg but also significantly reduces GPU memory usage during training to levels comparable to those during inference. Moreover, the proposed personalized FL strategy that is built upon the theoretical insights to customize the client-wise learning rate can effectively accelerate loss reduction. We hope our work can help to bridge theoretical and practical aspects of federated fine-tuning for LLMs, thereby stimulating further advancements and research in this area.
copyright: none
1. Introduction
Federated Learning (FL) has become an important approach in modern machine learning, particularly in scenarios where data decentralization and privacy-preserving are crucial (Muhammad et al., 2020; Meng et al., 2021; Hong et al., 2021; Wang et al., 2022; Xie et al., 2023). Central to this learning paradigm is the corroborative training of a global model through the aggregation of updates from multiple clients, without sharing their raw data (McMahan et al., 2017; Kairouz et al., 2021).
In parallel, Large Language Models (LLMs) have radically advanced the field of natural language processing (Touvron et al., 2023; Du et al., 2021; Brown et al., 2020). The fine-tuning of LLMs that are already pre-trained on vast corpora, has proven to be a highly effective strategy for numerous tasks, yielding models that are both versatile and capable of adapting to specific domain narratives or aligning with human values (Malladi et al., 2023; Brown et al., 2020).
The tuning of LLMs requires suitable alignment data, which are often costly to acquire (Conover et al., 2023b; Chen et al., 2024). Due to the abundance of private data that remains largely isolated and underutilized, the intersection of FL and LLMs has sparked increasing interest among researchers (Zhang et al., 2023a; Kuang et al., 2023; Fan et al., 2023a; Qin et al., 2024). Notably, this integration presents significant computational challenges, especially for clients with limited resources (Zhang et al., 2023b; Chen et al., 2023b). The scaling up of LLMs further compounds this issue, as the computation of gradients for backpropagation incurs substantial memory costs, frequently surpassing the practical capabilities of these clients (Malladi et al., 2023).
Addressing this challenge, we turn our attention to Zeroth-Order Optimization (ZOO), an algorithm that computes gradient approximations without explicit gradient information, thus significantly reducing memory consumption (Gu et al., 2021). However, the combination of ZOO and FL —a research direction we refer to as ZOO-FL —remains unexplored in the literature in the context of LLMs (Qiu et al., 2022). Our work intends to bridge this gap by harnessing the memory efficiency of ZOO within the context of federated fine-tuning of LLMs, especially on the following theoretical foundations:
(Q1) How does the vast parameter space of LLMs influence the behavior of ZOO-FL? (Q2) Can we establish the convergence properties of ZOO-FL for LLMs? (Q3) Which model parameters are critical for convergence, and how can we leverage them to optimize FL performance, such as via personalization?
In this paper, we focus on incorporating a memory-efficient ZOO method, MeZO (Malladi et al., 2023) into FL, a synergy we denote as FedMeZO, and establishing its convergence properties under the large-scale parameter space of LLMs. We analyze and present precise convergence rates characterized by the low effective rank $r$ of the models’ Hessian matrices (Aghajanyan et al., 2021; Li et al., 2018), and other typical FL parameters such as number of clients $N$, number of FL rounds $T$, iteration steps of local training $H$ and heterogeneity constants $c_{h}$ and $\sigma_{h}$. Moreover, we reveal the learning rate to be a crucial variable for convergence. Building on theoretical insights, we further propose a strategy that tailors the learning rate to each client’s specific data characteristics.
To validate our theoretical results, we conduct extensive experiments on real-world FL datasets for LLM tuning, which cover diverse data distributions and application tasks. Our empirical findings corroborate the theoretical analysis, validating effective convergence even when scaling up to models with billions of parameters. Compared with first-order methods such as FedAvg, FedMeZO converges faster meanwhile remarkably reducing GPU memory requirements. The personalized strategies guided by our theoretical insights, empirically show a more rapid loss reduction, as opposed to non-personalized or random learning rate assignments.
In summary, our theoretical and empirical exploration validates FedMeZO in the fine-tuning process of LLMs, providing a rigorous framework and practical insights for future applications. Our key contributions are threefold:
  • We advance the understanding of FedMeZO for LLMs, extending the two-point gradient estimation to federated tuning and establishing theoretical convergence rate $\mathcal{O}\left(r^{3/2}(NHT)^{-1/2}\right)$ and $\mathcal{O}\left(r^{3/2}(\widetilde{c}_{h}NHT)^{-1/2}\right)$ $-\mathcal{O}\Bigl{(}\sigma_{h}^{2}(c_{h}N)^{-1}\Bigr{)}$ for the i.i.d. setting and non-i.i.d. setting, respectively.
  • We analyze the impact of various hyperparameters of FedMeZO and explore a theory-informed strategy for personalized learning rate adjustment, providing practical guidance for ZOO-FL.
  • Through extensive empirical evidence with LLMs, we verify the proposed theoretical results and show that FedMeZO yields effective convergence with substantially reduced memory overhead compared to FedAvg. Our codes are publicly available at https://github.com/alibaba/FederatedScope/tree/FedMeZO.
2. Preliminaries
2.1. Background and Related Works
Federated Fine-Tuning of Large Language Models.
Large Language Models (LLMs) have demonstrated remarkable capabilities that enable a variety of real-world applications (Zhang et al., 2022; Touvron et al., 2023; Zhao et al., 2023). The federated fine-tuning of LLMs has recently attracted attention, focused on adapting these models to domain-specific tasks while preserving the privacy of the training data. Chen et al. (Chen et al., 2023a) investigated the integration of LLMs within federated settings, highlighting the inherent challenges and potential opportunities. Zhang et al. (Zhang et al., 2023b) furthered this research by examining instruction tuning of LLMs in a federated context, marking progress in applying FL to the specialized training of LLMs. Notable frameworks such as FATE-LLM by Fan et al. (Fan et al., 2023b) and FederatedScope-LLM by Kuang et al. (Kuang et al., 2023) offer industrial-grade and comprehensive solutions for federated fine-tuning. Our work, in contrast, investigates the fusion of Zeroth-Order Optimization (ZOO) with FL for the fine-tuning of LLMs, an area that has yet to be fully investigated, thereby addressing a gap in the literature and providing fundamental theoretical insights.
Zeroth-Order Optimization in Federated Learning.
ZOO has emerged as a viable method to address the difficulties of computing gradients in FL, especially in settings limited by computational resources. Zhang et al. (Zhang et al., 2021) proposed a ZOO algorithm tailored for vertical FL, focusing on privacy preservation. Yi et al. (Yi et al., 2022) and Li et al. (Li et al., 2021) studied ZOO-FL algorithms, with discussions on convergence properties with single-point perturbation and local updates in decentralized FL, respectively. The convergence analysis is a critical aspect of FL, as illustrated by Li et al. (Li et al., 2019) for the FedAvg algorithm and further developed by Fang et al. (Fang et al., 2022) for mini-batch stochastic ZOO-FL in wireless networks. Moreover, Shu et al. (Shu et al., 2023) proposed enhancements to query efficiency for ZOO within the FL framework. Our research sets itself apart by formulating theoretical convergence bounds for ZOO-FL, specifically tailored to the large-scale parameter space of LLMs. This builds on the preliminary work by Malladi et al. (Malladi et al., 2023), which confirmed the feasibility of ZOO for LLMs in a centralized setting.
2.2. Problem Formulation
For readability and brevity, we summarize the full list of introduced notations in Appendix A and present detailed proofs of all theoretical results in Appendix B-D.
Federated Learning.
We consider the general FL setting as of FedAvg (McMahan et al., 2017), with a central server and a collection of $N$ clients, indexed by ${1,2,...,N}$. The central server coordinates the training of a global model through the collaborative efforts of these clients, each holding local data samples drawn from their respective distributions $\mathcal{D}i$. The optimization problem can be formulated as:
(1) $\min_{\theta\in\mathbb{R}^{d}}f(\theta)\stackrel{{\scriptstyle\triangle}}{{=}}\frac{1}{N}\sum_{i=1}^{N}f_{i}(\theta_{i}),\quad f_{i}(\theta)\stackrel{{\scriptstyle\triangle}}{{=}}\mathbb{E}_{\mathcal{B}_{i}\sim\mathcal{D}_{i}}\bigl{[}F_{i}(\theta,\mathcal{B}_{i})\bigr{]},$
where $\theta\in\mathbb{R}^{d}$ denotes the $d$-dimension parameter of the model, and $f(\theta)$ and $f_{i}(\theta)$ denote the global loss function on the central server and local loss function on $i^{th}$ client, respectively. Typically, the clients are assumed with equal importance (Wang et al., 2021a), and the data is randomly sampled for efficiency (Li et al., 2014). $F_{i}(\theta,\mathcal{B}_{i})$ represents the local loss function w.r.t a specific mini-batch $\mathcal{B}_{i}$ drawn from $\mathcal{D}_{i}$.
Zeroth-Order Optimization.
Zeroth-order optimization (ZOO) is a prominent technique in scenarios where gradients are difficult to obtain, which estimates gradients by forward propagations. Given a random vector $z$ and a smoothing constant $\mu$, a typical one-point gradient estimator (Duchi et al., 2015) is defined as:
(2) ~$\widetilde{\nabla}F(\theta,z,\mathcal{B},\mu)=\frac{z}{2\mu}\bigl{(}F(\theta+\mu z,\mathcal{B})-F(\theta,\mathcal{B})\bigr{)},$
However, Eq. (2) provides a biased gradient estimation, leading to a certain degree of information loss (Liu et al., 2020). Hence our work employs the ZOO paradigm with a two-point gradient estimator proposed by (Malladi et al., 2023) in a federated setting:
Definition 2.1 ().
(Two-point gradient estimator) Given a set of parameters $\theta\in\mathbb{R}^{d}$ for an LLM and a mini-batch $\mathcal{B}_{i}$, the two-point zeroth-order gradient estimator is formulated as:
(3) ~$\widetilde{\nabla}F_{i}(\theta,z_{i},\mathcal{B}_{i},\mu)=\frac{z_{i}}{2\mu}\bigl{(}F_{i}(\theta+\mu z_{i},\mathcal{B}_{i})-F_{i}(\theta-\mu z_{i},\mathcal{B}_{i})\bigr{)},$
where $z_{i}\sim\mathcal{N}(0,I_{d})$ is a Gaussian random variable and $\mu$ is the perturbation scale. The two-point gradient estimator in Eq. (3) requires only two forward passes through the model to compute the estimation of gradient, which serves as a memory-efficient alternative to backpropagation (BP).
The FedMeZO Algorithm.
In this paper, we study and analyze the proprieties of a practical synergy of MeZO (Malladi et al., 2023) and FedAvg (McMahan et al., 2017), which is designed to fine-tune LLMs in an efficient, privacy-preserving and personalized manner. We term this ZOO-FL approach as FedMeZO, depicted with the following processes:
In a single communication round, the central server first broadcasts the global model parameters to available clients. Once the clients have completed their local updates and uploaded their models, the server aggregates the updates according to Eq. (1), forming the basis for the subsequent round.
Upon receiving the global model parameters, clients perform the following steps, distinguishing FedMeZO from traditional BP-based FedAvg algorithms in two-fold:
(1) Training Memory Reduction: Clients update their models using the two-point ZOO gradient estimator defined in Eq. (3) as:
(4) $e_{i}^{(t,k)}=\widetilde{\nabla}{F_{i}(\theta_{i}^{(t,k)},z_{i}^{(t,k)},\mathcal{B}_{i}^{(t,k)},\mu)},$
where $(t,k)$ denotes the $k^{th}$ iteration within the $t^{th}$ communication round. Unlike standard ZO-SGD algorithms that require storing the perturbation vector $z$ at each iteration, FedMeZO resamples $z$ using random seeds in in-place implementation, thus reducing memory usage to a level equivalent to inference (Malladi et al., 2023).
(2) Communication Cost Reduction: To mitigate the high communication overhead associated with LLMs, FedMeZO leverages Low-Rank Adaptation (LoRA) (Hu et al., 2022), which introduces reparametrization to tune two small delta matrix on the linear layers instead of the whole LLM weights, based on the assumption that well pre-trained LLM possess a low “intrinsic dimension” when adapted to new tasks. Introducing it can help us further reduce the number of parameters to be updated and uploaded, thereby aligning with the practical constraints of federated settings.
2.3. Lemmas and Assumptions
Lemma 2.2.
(Unbiased Gradient Estimator) The two-point zeroth-order gradient estimator described in Eq. (3) is an unbiased estimator of the true gradient, that is,
(5) $\mathbb{E}[\widetilde{\nabla}F_{i}(\theta,z_{i},\mathcal{B}_{i},\mu)]=\nabla f_{i}(\theta).$
The Hessian matrix, which is the square matrix of second-order partial derivatives of the loss w.r.t the model parameters, characterizes the curvature of the loss surface (Ghorbani et al., 2019). Although the size of a model’s loss Hessian is often associated with the rate of fine-tuning, studies suggest that the large-scale parameters of LLMs do not necessarily impede convergence (Jamieson et al., 2012; Agarwal et al., 2009). This paradox is addressed by recognizing that the loss Hessian often exhibits a small local effective rank (Malladi et al., 2023), which we capture in the following assumption:
Assumption 1.
There exist a Hessian matrix $\mathcal{H}(\theta^{t})$ satisfying:
$\bullet$ $\nabla^{2}f(\theta)\preceq\mathcal{H}(\theta^{t})$ for all $\theta$ such that $\lVert\theta-\theta^{t}\rVert\leq\eta dG(\theta^{t})$, where $G(\theta^{t})=\max_{\mathcal{B}\sim\mathcal{D}}\lVert\nabla f(\theta^{t},\mathcal{B})\rVert$.
[…]
Assumption 1 characterizes a low effective rank $r$ in the Hessian matrix, which demonstrates that LLM fine-tuning can occur in a low dimensional subspace ($\leq 200$ parameters) (Aghajanyan et al., 2021; Li et al., 2018). With this insight, (Malladi et al., 2023) identified the bound of loss descent at each step of centralized ZOO, which is partially influenced by $r$:
Lemma 2.3.
(Bounded Centralized Descent) Assume $f(\theta)$ is $L$-smooth and let ~$\widetilde{\nabla}{F(\theta,z,\mathcal{B},\mu)}$ be the unbiased zeroth-order gradient estimator from Eq. (3). If the Hessian matrix $\mathcal{H}(\theta)$ exhibits a local effective rank of $r$, and constants $\gamma=\Theta(r/n)$ and $\zeta=\Theta(1/rd)$ exist, then the expected decrease in loss can be bounded as follows:
$\mathbb{E}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})$ $-\frac{\eta}{\gamma}{\lVert\nabla f(\theta^{t})\rVert}^{2}$
(6) $+\frac{\eta^{2}L\zeta}{2}\mathbb{E}\bigl{[}{\lVert\widetilde{\nabla}{F(\theta,z,\mathcal{B},\mu)}\rVert}^{2}\bigr{]},$
where $\gamma=\frac{dr+d-2}{n(d+2)}$, $\zeta=\frac{(d+2)n^{2}}{(dr+d-2)(d+n-1)}$, and $n$ is the number of randomizations.
From Eq. (2.3), we observe that the rate of descent at a single step depends on the gradient related to $\gamma$ and the gradient estimation related to $\zeta$. Following (Malladi et al., 2023), we set $n$ to 1 in this paper.
Besides, to facilitate the analysis in FL setting, we introduce four assumptions, including Bounded Loss (Assumption 2), $L$-smoothness (Assumption 3), mini-batch gradient error bound (Assumption 4, global-local disparities in i.i.d. and non-i.i.d. settings (Assumptions 5 and 6 respectively). These assumptions are standard and foundational in optimization and FL literature (Bottou et al., 2018; Li et al., 2019; Wang et al., 2021b; Li et al., 2020), which we detail in Appendix B.
3. Main results
3.1. Convergence Analysis in i.i.d. Case
In this subsection, we examine the convergence properties of FedMeZO within the i.i.d. data distribution setting. We establish the conditions under which the algorithm guarantees loss reduction at each iteration and provide a global convergence rate.
Theorem 3.1.
(Stepwise Loss Descent in i.i.d. Setting) Under Assumptions 1-5 and with a learning rate $\eta$ satisfying
(7) $\eta\leq\min\Bigl{\{}\frac{1}{3HL\sqrt{c_{g}d}},\frac{N}{3HLc_{g}},\frac{1}{H^{2}}\Bigr{\}},$
the expected decrease in loss at each step for FedMeZO under the i.i.d. scenario is bounded as
$\mathbb{E}_{t}\left[f(\theta^{t+1})\right]$ $\leq f(\theta^{t})-\left(\frac{2}{\gamma}-\frac{2\zeta}{d}\right)\eta\mathbb{E}_{t}\left\lVert\nabla f(\theta^{t})\right\rVert^{2}$
(8) $\quad+\frac{2\sigma_{g}^{2}\zeta\eta L}{NHd}+\frac{\zeta\eta\mu^{2}L^{3}}{2NH},$
where $\gamma$ and $\zeta$ quantify the effective low-rank properties of the gradient and its estimator, respectively.
In Theorem 3.1, the term $(-2\eta/\gamma)\mathbb{E}_{t}|\nabla f(\theta^{t})|^{2}$ serves as a critical factor that drives the decrease in the loss function, as it is the sole negative contributor in Eq. (3.1) such that $\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})-f(\theta^{t})\bigr{)}\bigr{]}\leq 0$. Note that the presence of the factor $\gamma^{-1}=\Theta(r^{-1})$ underscores the impact of the low effective rank $r$ on the convergence rate (under Assumption 1), revealing that a reduction in $r$ can accelerate convergence independently of the high-dimensional parameter space $d$. Consequently, even for LLMs with expansive parameter spaces, FedMeZO can attain convergence. This addresses our first foundational question “Q1: How does the vast parameter space of LLMs influence the behavior of ZOO-FL?”.
Moreover, the terms $(2\eta\zeta/d)\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$ and $(2\sigma_{g}^{2}\zeta\eta L/NHd)$ are scaled by $\zeta/d$, i.e., in $\Theta(1/rd^{2})$, contributing a relatively smaller effect on the convergence speed compared to negative term. This demonstrates that the influence on convergence speed from the zeroth-order gradient estimation is moderated by the model’s effective low rank and dimensionality. As for the last term, $(\zeta\eta\mu^{2}L^{3}/2NH)$, it acts as a factor slowing down the convergence rate, and we can observe that when $N$ and $H$ are larger, this term becomes smaller. This suggests that the effect of slowing down the convergence rate is not as pronounced, and simultaneously, the perturbation step $\mu$ should not be excessively large. Specifically, this term indicates that increasing the number of clients and the number of local rounds can enhance convergence, while also emphasizing the importance of keeping the perturbation step $\mu$ moderate.
After gaining intuitive insights in each round of training through the analysis of Theorem 3.1, it is necessary to assess the convergence performance of FedMeZO from a global perspective. We utilize the squared magnitude of the gradient $\mathbb{E}_{t}\lVert\nabla f(\theta^{t})\rVert^{2}$ as a measure to assess the suboptimality of each iterate. The rapidity with which the algorithm approaches a stationary point serves as a crucial metric for determining its efficacy in the context of non-convex optimization problems (Nesterov and Spokoiny, 2017).
Corollary 3.2.
(Global Convergence in i.i.d. Setting) Assuming the conditions of Theorem 3.1 hold, the global convergence for FedMeZO in the i.i.d. case, characterized by $\Gamma=\frac{d-\zeta\gamma}{d\gamma}$, is given by
(9) $\min_{t\in[T]}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq\frac{f(\theta^{0})-f^{*}}{2\eta T\Gamma}+\frac{\sigma_{g}^{2}\zeta L}{NHd\Gamma}+\frac{\zeta\mu^{2}L^{3}}{4NH\Gamma}$
where $f^{*}$ denotes the optimal loss value.
The upper bound on the minimum squared gradient norm across iterations is composed of three terms in Corollary 3.2. The first term indicates that the distance to the optimal loss relies on the initial state of optimality, while the second and third terms elucidate the influences of stochastic mini-batch errors and the perturbation scale $\mu$ inherent to ZOO, respectively. Specifically, they both reflect the impact of the model parameters $d$ and the low effective rank $r$ on the optimal loss. As pointed out in (Nesterov and Spokoiny, 2017), by choosing an appropriate step size, we can obtain the desired accuracy.
Given that $\gamma=\Theta(r)$ and $\zeta=\Theta\left(\frac{1}{rd}\right)$, we have $\zeta\gamma=\Theta\left(\frac{r}{rd}\right)=\Theta\left(\frac{1}{d}\right)$. As the parameter $d$ is large, $d-\zeta\gamma=d-\Theta\left(\frac{1}{d}\right)$ is dominated by $d$. Consequently, the dominant term of $\Gamma$ is $\frac{d\gamma}{d-\zeta\gamma}$, which simplifies to $\gamma=\Theta(r)$. Therefore, $\Gamma=\Theta\left(\frac{1}{\gamma}\right)=\Theta\left(\frac{1}{r}\right)$. Building on this relationship, we have the following corollary that articulates the convergence rate of FedMeZO.
Corollary 3.3.
(Convergence Rate in i.i.d. Setting) Assuming the conditions of Corollary 3.2 hold and given $\eta=(NH)^{1/2}(rT)^{-1/2}$ and $\mu=(NH)^{1/4}r^{-1/2}$, we have
(10) $\min_{t\in[T]}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq\mathcal{O}\left(r^{3/2}(NHT)^{-1/2}\right)+\mathcal{O}\left(d^{-1}(rNH)^{-1/2}\right).$
The expression on the right-hand side of Eq. (10) is dominated by $\mathcal{O}\left(r^{3/2}(NHT)^{-1/2}\right)$. Consequently, we have derived the convergence rate for FedMeZO. The low effective rank $r$ significantly contributes to lowering the convergence rate, which is also influenced by the number of clients $N$, the steps of local training iteration $H$, and the total number of communication rounds $T$. Moreover, to satisfy the learning rate condition in Eq. (7), the values of $N$, $H$, and $T$ must be suitably large.
It is important to note that FedMeZO does not primarily aim to accelerate convergence speed but rather to identify the convergence rate under assumptions pertinent to LLMs. This is intended to demonstrate that FedMeZO can achieve convergence even within a vast parameter space. In a series of studies on federated ZOO, Federated Zeroth-Order Optimization (FedZO) presents the most comprehensive and complete analysis with a convergence rate of $\mathcal{O}\left(\sqrt{d/(NHTb_{1}b_{2})}\right)$ (Fang et al., 2022), which exhibits a lower rate compared to $\mathcal{O}\left(d^{3}/T\right)$ of ZONE-S (Hajinezhad et al., 2019) and accounts for the impact of $H$ compared to $\mathcal{O}\left(\sqrt{d/NT}\right)$ of DZOPA (Yi et al., 2022). In contrast to FedZO, our method, FedMeZO, theoretically supports a faster convergence by replacing $d^{1/2}$ with $r^{3/2}$ and setting $b_{1}=b_{2}=1$. These comparisons show that FedMeZO addresses the challenges posed by large models, offering an efficient convergence rate that relies on $r$.
This advancement signifies progress in optimizing federated learning algorithms, particularly for LLMs, where the scalability of parameters and data heterogeneity are major challenges. By emphasizing the low effective rank, our approach enhances both the theoretical understanding of convergence behavior in complex settings and the guidance insights into the settings of learning rates and other parameters to achieve efficient convergence outcomes.
However, i.i.d. data is typically encountered in idealized environments. In real-world applications, non-i.i.d. conditions are more common and challenging. Next, we further discuss and analyze the convergence of FedMeZO under non-i.i.d. settings.
3.2. Convergence Analysis in Non-i.i.d Case
Analyzing convergence in the context of non-i.i.d. data distributions is crucial for understanding the behavior of FL algorithms in real-world scenarios. In this section, we extend our convergence analysis to the case where client data distributions are heterogeneous.
Theorem 3.4.
(Stepwise Loss Descent in Non-i.i.d Setting) Let Assumptions 1-4 and Assumption 6 hold and learning rate $\eta$ satisfy
(11) $\eta\leq\min\Bigl{\{}\frac{1}{3HL\sqrt{c_{g}d}},\frac{N}{3HLc_{g}},\frac{1}{H^{2}}\Bigr{\}}.$
Then, the expected loss at each step for FedMeZO in the non-i.i.d. setting is bounded as
$\mathbb{E}_{t}\left[f(\theta^{t+1})\right]$ $\leq f(\theta^{t})-\left(\frac{2}{\gamma N}-\frac{2\zeta\widetilde{c}_{h}}{d}\right)\eta\mathbb{E}_{t}\left\lVert\nabla f(\theta^{t})\right\rVert^{2}$
(12) $\quad+\frac{2\widetilde{\sigma}^{2}\zeta L\eta}{NHd}+\frac{\zeta\eta\mu^{2}L^{3}}{2NH}-\frac{2}{\gamma N}\eta\sigma_{h}^{2},$
where $\widetilde{c}_{h}=c_{h}+N$ and $\widetilde{\sigma}^{2}=3c_{g}\sigma_{h}^{2}+\sigma_{g}^{2}$.
Comparing Eq. (3.4) with its i.i.d. counterpart Eq. (3.1), the non-i.i.d. setting introduces additional terms reflecting data heterogeneity. Firstly, an additional term $\widetilde{c}_{h}$ appears before $\mathbb{E}_{t}\lVert\nabla f(\theta^{t})\rVert^{2}$; secondly, the original $\sigma_{g}^{2}$ change into $\widetilde{\sigma}^{2}$; thirdly, a new term related to $\sigma_{h}^{2}$ is added at the end. The term $\widetilde{c}_{h}$ amplifies the effect of the gradient norm, while $\widetilde{\sigma}^{2}$ encapsulates both the intrinsic stochasticity and data heterogeneity. The presence of $\sigma_{h}^{2}$ indicates the impact of client data divergence on the convergence behavior, in a degree dependent on $\Theta(1/rd^{2})$. Given that the contribution of the negative term accelerates the rate of decline in each round, it can be concluded that heterogeneity is positively correlated with convergence. Considering all the above changes, appropriate heterogeneity can aid in the model convergence.
Experimental results in Section 5.3.3 confirm that a more randomized dataset distribution leads to improved convergence, supporting our theoretical insights. Next, we present the global convergence result for the non-i.i.d. setting building upon Theorem 3.4.
Corollary 3.5.
(Global Convergence in Non-i.i.d. Setting) Assuming the conditions of Theorem 3.4 hold, denote $\widetilde{\Gamma}=\frac{d-N\gamma\zeta}{d\gamma N}$, FedMeZO satisfies:
$\min_{t\in[T]}\mathbb{E}_{t}\left\lVert\nabla f(\theta^{t})\right\rVert^{2}$ $\leq\frac{f(\theta^{0})-f^{*}}{2\widetilde{\Gamma}\widetilde{c}_{h}\eta T}+\frac{\widetilde{\sigma}^{2}\zeta L}{\widetilde{\Gamma}\widetilde{c}_{h}NHd}$
(13) $\quad+\frac{\zeta\mu^{2}L^{3}}{4\widetilde{\Gamma}\widetilde{c}_{h}NH}-\frac{\sigma_{h}^{2}}{\widetilde{\Gamma}\widetilde{c}_{h}\gamma N}.$
Given $\gamma=\Theta(r)$ and $\zeta=\Theta\left(\frac{1}{rd}\right)$, the expression $d-N\zeta\gamma$ simplifies to $d-\Theta\left(\frac{N}{d}\right)$ which is dominated by $d$. Consequently, $\widetilde{\Gamma}$ simplifies to $\Theta\left(\frac{1}{r}\right)$ as in the non-i.i.d. case. Compared to Corollary 3.2, Corollary 3.5 introduces two changes: first, all terms on the right side of Eq. 3.5 include a denominator $c_{g}$, and second, there is an additional term associated with non-i.i.d. heterogeneity, scaled with $\Theta(\sigma_{h}^{2}/c_{h}N)$. This further demonstrates the constraining effect of data heterogeneity in FedMeZO. Similar to Corollary 3.3, by setting appropriate values for $\eta$ and $\mu$, we obtain the following convergence rate.
Corollary 3.6.
(Convergence Rate in Non-i.i.d. Setting) Assuming the conditions of Corollary 3.5 hold, with $\eta=(NH)^{1/2}(r\widetilde{c}_{h}T)^{-1/2}$ and $\mu=(\widetilde{c}hNH)^{1/4}r^{-1/2}$, FedMeZO has convergence rate as follows:
$\min_{t\in[T]}\mathbb{E}_{t}\Big{\lVert}\nabla$ $f(\theta^{t})\Big{\rVert}^{2}\leq\mathcal{O}\left(r^{3/2}(\widetilde{c}_{h}NHT)^{-1/2}\right)$
(14) $\quad+\mathcal{O}\left(d^{-1}(r\widetilde{c}_{h}NH)^{-1/2}\right)-\mathcal{O}\left(\sigma_{h}^{2}(c_{h}N)^{-1}\right).$
The convergence rate in Eq. (3.6) is primarily driven by the term $\mathcal{O}\left(r^{3/2}(\widetilde{c}_{h}NHT)^{-1/2}\right)$, indicating that optimizing the balance between $\widetilde{c}_{h}$, $c_{h}$, and $N$ is crucial, which reflects a complex interplay of heterogeneity. Specifically, we observe that to achieve better convergence, a smaller $\mathcal{O}\Bigl{(}r^{\frac{3}{2}}(\widetilde{c}_{h}NHT)^{-\frac{1}{2}}\Bigr{)}$ is preferred while the $\mathcal{O}\Bigl{(}\sigma_{h}^{2}(c_{h}N)^{-1}\Bigr{)}$ term need to increase at the same time. Consequently, the balance between $c_{g}$, $c_{h}$, and $N$ becomes a dynamic trade-off process, i.e., the heterogeneity among different clients directly influences the overall convergence performance.
For now, we have answered the question “Q2: Can we establish the convergence properties of ZOO-FL for LLMs?” via theorems and corollaries mentioned in this section. We also validated the nature of convergence under different scenarios and tasks through empirical experiments in Section 5.2.
3.3. Implications
The aforementioned theoretical results offer numerous insights into parameter tuning. A critical revelation from our analysis pertains to the constraints imposed on the learning rate, as delineated in Eq. (7) and Eq. (11), which suggests that an optimal learning rate magnitude is anchored at $1/\sqrt{d}$. Larger learning rates are not only ineffectual but also pose a risk of destabilizing the training dynamic. In Appendix F.1, our empirical experiments corroborate this hypothesis, demonstrating that excessive learning rates precipitate abrupt increases in loss.
Furthermore, our insights regarding the learning rate open up prospects for personalized FL, a compelling approach that uses client-specific configurations to address heterogeneity and has attracted increasing interest (Sattler et al., 2020; Chen et al., 2022; Fallah et al., 2020; Marfoq et al., 2020; Chen et al., 2023c). Specifically, we investigate theory-guided personalized strategies by dynamically adjusting the learning rate $\eta_{i}$ in proportion to a quantifiable measure of data heterogeneity among clients. In light of Theorem 3.4 that a larger heterogeneity is more conducive to model convergence, it is feasible to appropriately increase the learning rate, allowing specific clients to contribute more to the overall convergence, we propose the following tailored adjustment strategy:
Proposition 3.7.
(Adaptive Learning Rate Adjustment) Let Assumption 6 hold, the learning rate $\eta_{i}$ can be adjusted according to the formula to better accommodate the varied learning landscapes than non-personalized FL:
(15) $\eta_{i}=\eta_{0}(1+\alpha\cdot\Phi_{i}),$
where $\eta_{0}$ represents a default learning rate applicable in a i.i.d. setting, $\alpha$ is a scaling factor that determines the sensitivity of the learning rate and $\Phi_{i}$ is the heterogeneity index, representing the extent of $c_{g}$ and $\sigma_{h}^{2}$. This proposition underscores the importance of considering data heterogeneity in the design of the learning rate strategy within personalized FL, offering a structured approach to enhance learning outcomes across diverse client datasets.
In Section 5.4, we empirically confirm that a particular implementation of this strategy facilitates faster convergence. It is important to note that the data heterogeneity index $\Phi_{i}$ cannot be determined a priori; therefore, we utilize several proxy measures during the training process to estimate it. Our goal is not to prescribe an exact solution to this strategy, but rather, through analysis and empirical investigation, to enlighten further research and development of personalized FedMeZO for more effective training of LLMs.
These discussions and corresponding empirical support address the question “Q3: Which model parameters are critical for convergence, and how can we leverage them to optimize FL performance, such as via personalization?”.
Besides, recall that we adopt LoRA to mitigate the communication burden associated with LLMs for practical FL scenarios. Nonetheless, the influence of LoRA on the model’s low effective rank, remains an open question. We thus advance the following conjecture under Assumption 1, predicated on existing literature(Yao et al., 2020; Sagun et al., 2017), to facilitate further validations:
Conjecture 3.8 (Rank Correlation).
[…]
4. Proof Outline
This section provides an outline of the derivations presented in Section 3, emphasizing the key analytical techniques and concepts employed. Detailed proofs are available in Appendix D.
We begin by taking expectations on both sides of Eq. (2.3), considering a federated learning setting. The equation is split into two main parts for further analysis:
$\small\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})-\frac{1}{\gamma}\cdot\eta\mathbb{E}_{t}\Big{\lVert}\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}(\theta^{t})\Big{\rVert}^{2}$
(16) $+\frac{1}{2}\eta^{2}L\cdot\zeta\cdot\mathbb{E}_{t}\Big{\lVert}\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}e_{i}^{(t,k)}\Big{\rVert}^{2}.$
For simplicity, we denote the two expectation terms as $T_{1}$ and $T_{2}$, which pertain to the expected squared norms of the gradient and the zeroth-order gradient estimator, respectively.
4.1. Proof of Theorem 3.1
For term $T_{1}$ in Eq. (4), we utilize the Cauchy-Schwarz inequality ${\lVert a+b\rVert}^{2}\leq 2{\lVert a\rVert}^{2}+2{\lVert b\rVert}^{2}$ to decompose it into two parts, with the first representing the discrepancy between local and global gradients. By invoking Assumption 5, we establish:
$T_{1}\leq\frac{2}{N}\sum_{i=1}^{N}\mathbb{E}_{t}{\Big{\lVert}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})\Big{\rVert}}^{2}+2\cdot\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}.$
For term $T_{2}$, Jensen’s inequality allows us to bound the expected squared norm of the zeroth-order gradient estimator as follows:
(17) $T_{2}\leq\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}e_{i}^{(t,k)}\Big{\rVert}^{2}.$
By substituting Eq. (3) into Eq. (17), we proceed to use the Cauchy-Schwarz inequality to decompose this gradient estimator into two parts, each of which is a biased estimator. Recall that in our gradient estimator, $z_{i}$ follows a Gaussian distribution. Therefore, the impact on the norm caused by a forward step and a backward step of the estimator is identical. Consequently, we ascertain that the term $T_{2}$ is bounded by a single-point gradient estimation as $\mathbb{E}_{t}\Bigg{\lVert}\frac{z_{i}^{(t,k)}}{\mu}\Bigl{(}F_{i}(\theta_{i}^{(t,k)}+\mu z_{i}^{(t,k)},\mathcal{B}_{i}^{(t,k)})$ $-F_{i}(\theta_{i}^{(t,k)},\mathcal{B}_{i}^{(t,k)})\Bigr{)}\Bigg{\rVert}^{2}$.
Following Lemma 4.1 in (Gao et al., 2018), we can bound the expectation term as:
$\mathbb{E}_{t}\Big{\lVert}e_{i}^{(t,k)}$ $\Big{\rVert}^{2}\leq\frac{1}{d^{2}}\Biggl{[}2d\cdot\mathbb{E}_{t}\Big{\lVert}\nabla F_{i}(\theta_{i}^{(t,k)},\mathcal{B}_{i}^{(t,k)})\Big{\rVert}^{2}+\frac{\mu^{2}}{2}L^{2}d^{2}\Biggr{]}$
(18) $\leq\frac{1}{d^{2}}\Biggl{[}2c_{g}d\cdot\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}+2d\sigma_{g}^{2}+\frac{\mu^{2}}{2}L^{2}d^{2}\Biggr{]},$
where the second inequality is derived based on Assumption 4. Subsequently, we bound the expectation term by applying the Cauchy-Schwartz inequality to divide it into three parts:
$\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}$ $=\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\mp\nabla f_{i}(\theta^{t})\mp\nabla f(\theta_{i}^{t})\Big{\rVert}^{2}$
(19) $\leq 3L^{2}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+3\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}.$
The first part represents the gradient difference between stages $(t,k)$ and $(t,0)$, which can be computed using Assumption 3, i.e., the $L$-smooth condition. The second part signifies the disparity between local and global aspects, calculated using Assumption 5. The third part is retained as is.
Combining Equations (17), (4.1) and (4.1), we bound $T_{2}$ as follow:
$T_{2}\leq\frac{6c_{g}L^{2}}{Nd}$ $\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$
(20) $+$ $\frac{6c_{g}H}{d}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+\frac{2H\sigma_{g}^{2}}{d}+\frac{\mu^{2}HL^{2}}{2}.$
Next, combining Equations (4), (4.1) and (4.1), we have:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}-f(\theta^{t})$ $\leq\Biggl{(}\frac{3c_{g}\zeta\eta^{2}HL}{d}-\frac{2\eta}{\gamma}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$+\frac{3c_{g}\zeta\eta^{2}L^{3}}{Nd}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$
(21) $+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{d}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4}.$
In Eq. (4.1), $\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$ remains unknown and we need to constrain it further. The key idea is to transform this expectation term into a form related to $\mathbb{E}_{t}^{k}\lVert e_{i}^{(t,k)}\rVert^{2}$ and then utilize the conclusion of Eq. (4.1) and Eq. (4.1) for computation. The detailed derivation process is provided in the Appendix D.1 and we can have the bounded result:
(22) $\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}\leq\frac{C_{1}}{C_{0}},$
where $C_{0}=1-3c_{g}d\eta^{2}H^{2}L^{2}$ and $C_{1}=2c_{g}dH^{3}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$ $+\frac{2}{3}d\sigma_{g}^{2}H^{3}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}H^{3}\eta^{2}}{6}$.
Finally, by substituting Eq. (22) into Eq. (4.1), we obtain the final result of the stepwise descent. After simplifying and appropriately setting the learning rate, we arrive at the result presented in Theorem 3.1. The detailed derivation of this result is provided in Appendix D.1.
4.2. Proof of Theorem 3.4
The proof for the non-i.i.d. case follows a similar structure to that of the i.i.d. case, with adjustments made for the heterogeneity between local and global models as captured by $c_{h}$ and $\sigma_{h}^{2}$ (Assumption 6). In particular, we redefine $T_{1}$ in Eq. (4) as $\widetilde{T}_{1}$ to reflect the increased variance due to non-i.i.d. data:
$\widetilde{T}_{1}$ $\leq\frac{2}{N}\sum_{i=1}^{N}\mathbb{E}_{t}{\Big{\lVert}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})\Big{\rVert}}^{2}+2\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}$
(23) $\leq 2(1+c_{h})\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}+2\sigma_{h}^{2},$
where the second inequality follows Assumption 6.
Subsequently, $\widetilde{T}_{2}$ is computed similarly, with the heterogeneity terms incorporated. The difference from the i.i.d. case lies in the bounding of expectation term in Eq. (4.1):
$\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}$ $\leq 3L^{2}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$
(24) $\qquad+3(c_{h}+1)\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+3\sigma_{h}^{2},$
where the inequality employs Assumption 6.
Combining Equations (17), (4.1) and (4.2), we bound $\widetilde{T}_{2}$ as follow:
$\widetilde{T}_{2}\leq\frac{6c_{g}L^{2}}{Nd}\sum_{i=1}^{N}\sum_{k=1}^{H}$ $\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+\frac{6c_{g}(c_{h}+1)H}{d}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
(25) $+\frac{6c_{g}\sigma_{h}^{2}H}{d}+\frac{2H\sigma_{g}^{2}}{d}+\frac{\mu^{2}HL^{2}}{2}.$
By substituting $\widetilde{T}_{1}$ in Eq. (4.2) and $\widetilde{T}_{2}$ in Eq. (4.2) into Eq. (4), we get the result under the non-i.i.d. condition as follow:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}-f(\theta^{t})$ $\leq\Biggl{(}\frac{3c_{g}\widetilde{c}_{h}\zeta\eta^{2}HL}{d}-\frac{2\widetilde{c}_{h}\eta}{\gamma}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$+\frac{3c_{g}\zeta\eta^{2}L^{3}}{Nd}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$
(26) $+\frac{\widetilde{\sigma}^{2}\zeta\eta^{2}HL}{d}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4}-\frac{2}{\gamma}\sigma_{h}^{2}\eta,$
where $\widetilde{c}_{h}$ denotes $(c_{h}+1)$ and $\widetilde{\sigma}^{2}$ denotes $(3c_{g}\sigma_{h}^{2}+\sigma_{g}^{2})$. We then need to make the expectation term bounded. Unlike Eq. (22), due to the variations introduced by Eq. (4.2), two additional terms related to $\widetilde{c}_{h}$ and $\widetilde{\sigma}^{2}$ emerge, yielding the following result:
(27) $\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-$ $\theta^{t}\Big{\rVert}^{2}\leq\frac{C_{2}}{C_{0}},$
where $C_{1}=1-3c_{g}d\eta^{2}H^{2}L^{2}$ and $C_{2}=2c_{g}d\widetilde{c}_{h}H^{3}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$ $+\frac{2}{3}\widetilde{\sigma}^{2}dH^{3}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}H^{3}\eta^{2}}{6}$.
Finally, we can substitute Eq. (27) into Eq. (4.2) and have the result of Theorem 3.4. The detailed derivations about these steps of Theorem 3.4 are provided in Appendix D.3.
5. Empirical Support
This section aims to empirically validate our theoretical findings through a series of experiments.
5.1. Experimental Setup
We utilize LLaMA-3B (Touvron et al., 2023) as the foundational model and employ four datasets covering a range of tasks and data distribution types to provide comprehensive validation of our theoretical results (Kuang et al., 2023). Given that our theory centers on the loss function, we primarily focus on analyzing loss descent in our experiments. More details of the used datasets are in the Appendix (Table 3).
We set the total number of communication rounds to 500. By default, BP-based baselines undertake local training for one epoch, whereas FedMeZO conducts local training for 30 steps. We repeat our experiments with three seeds and plot the error bars. For more detailed implementation specifics, please refer to Appendix E.
5.2. Convergence Study
To assess the convergence of FedMeZO, we perform experiments on three datasets using different data splitters, as specified in Table 3, with test loss serving as the convergence metric. Our objective is to evaluate the generalization and stability of FedMeZO across diverse datasets and heterogeneity scenarios. For benchmarking purposes, we also measure the performance of BP-based FedAvg on the same datasets. Additionally, we document the GPU memory usage during training in Table 1. Representative findings are illustrated in Figure 1, while all the comprehensive results are available in Appendix F.6 due to the space limitation.
Figure 1. Convergence comparison of FedMeZO and BP-based FedAvg. More results are in Appendix F.6.
Table 1. The GPU Memory of BP-based FedAvg and FedMeZO.
Task BP-based FedAvg (MiB) FedMeZO (MiB)
Dolly-Meta 26571 10061
GSM8K-IID 17771 9733
CodeAlpaca-LDA 15287 9569
Two main conclusions emerge from the convergence experiments: First, when the learning rate complies with the requirements discussed in Section 3.3, as stipulated by Theorem 3.1 and Theorem 3.4, FedMeZO consistently diminishes loss with each step, ultimately achieving stable convergence. Second, under equivalent learning rate configurations, FedMeZO decreases loss more rapidly than BP-based FedAvg, indicating a swifter convergence rate. For instance, in the Dolly-Meta figure, FedMeZO stabilizes and converges around 300 rounds, whereas BP-based FedAvg’s loss is still declining at this juncture. Notably, from Table 1, we observe that the GPU memory demand for FedMeZO is roughly one-third of that required by BP-based FedAvg, suggesting that FedMeZO can achieve a speedier convergence with fewer resources.
5.3. Hyper-parameters Study
In this subsection, we perform a series of experiments to ascertain the influence of various hyper-parameters, as intimated by our theoretical findings.
5.3.1. Impact of Perturbation Scale $\mu$
To corroborate the theoretical impacts of the perturbation step on convergence, we examine $\mu$ values of $5\times 10^{-3}$ and $2\times 10^{-4}$, in addition to the default $\mu=1\times 10^{-3}$. We leverage the same datasets and splitters as in Section 5.2 for robustness. Figure 2 shows representative outcomes, with comprehensive results in Appendix F.4.
Figure 2. Effects of different perturbation scales $\mu$. More results are in Appendix F.4.
The results confirm that, consistent with Equations (10) and (3.6), a smaller $\mu$ marginally expedites model convergence. Figure 2 exemplifies that the training trajectory with $\mu=2\times 10^{-4}$ descends more rapidly than the others. However, given that $\mu$ appears as a second-order term in $\frac{\zeta\mu^{2}L^{3}}{4\widetilde{\Gamma}\widetilde{c}_{h}NH}$ and its absolute value is relatively small, its overall influence is modest. This is evident in Figure 2, where modifications to $\mu$ within a specific range yield only slight variations. Thus, a smaller $\mu$ proves advantageous for model convergence.
5.3.2. Impact of Local Iteration Step $H$
To validate the theoretical impact of local iteration steps on convergence, we contrast $H=10$ and $H=50$ with the standard $H=30$. Utilizing identical datasets and splitters from Section 5.2, we present typical findings in Figure 3, with all the detailed results forthcoming in Appendix F.5.
Figure 3. Effects of different local iteration steps $H$. More results are in Appendix F.5.
These experimental results suggest that a lower $H$ engenders a more sluggish convergence pace, whereas a higher $H$ somewhat propels convergence, mirroring the impact of $H$ as a denominator in the theoretical convergence rate analysis. Nonetheless, an excessive $H$ may lead to instability, as depicted by the curve of $H=50$, which exhibits a surge endwise in the figure. Hence, an appropriate choice of $H$ facilitates efficient model convergence.
5.3.3. Analysis of Other Hyper-parameters
We also explore the ramifications of learning rate, data splitters and the number of clients $N$ on the convergence rate. Due to the space limit, details pertaining to these experimental settings and results are presented in Appendices F.1, F.2 and F.3, respectively.
In a nutshell, as shown in Fig. 5, a suitable learning rate anchored at $1/\sqrt{d}$ leads to stabilized training dynamics, while larger ones exhibit divergence as suggested by our analysis (Section 3.3). Besides, dissimilar splitters symbolize varying extents of data heterogeneity, and we have observations revealing that augmented heterogeneity culminates in lower stabilized loss values (as shown in Fig. 6). This intimates that a moderate degree of data heterogeneity can elevate the model’s convergence proficiency. Moreover, Fig. 7 verifies our theoretical analysis in Section 3 that an increase in $N$ helps stabilize the global convergence.
5.4. Personalization Study
To reconcile Proposition 3.7 with practical scenarios, we conduct the following subsequent experiments. To account for each client’s heterogeneity during model updates in each round, we derive three signal quantities: (1) Round-wise Train Loss Difference: The discrepancy between each client’s loss in the preceding training round and the global loss. (2) Five-round Average Train Loss Difference: The average loss deviation for each client relative to the global loss over the antecedent five rounds. (3) Model Parameter Update Difference: The disparity between each client’s previous round parameter updates and the global update magnitude. We normalize them to the range of $(-1,1)$, serving as empirical estimates for $\Phi$.
For the setting of the scaling factor $\alpha$, following the guidance of the learning rate in Section 3.3, we designate $1.5\times 10^{-5}$ as the maximal learning rate, potentially leading to surges as per the learning rate search network. Symmetrically, we posit the minimal value at $5\times 10^{-6}$, anchored on the default learning rate of $1\times 10^{-5}$, thereby assigning $\alpha$ a value of $5\times 10^{-6}$.
Figure 4. Comparison of different strategies of learning rate adjustment. “Default” indicates non-personalized case, and “Round-wise Loss”, “Five-round Loss” and “Model Update Difference” indicate three signal quantities leveraged.
As counterpoints, we furnish two configuration strategies for the learning rate adjustment: one uniformly applies a default learning rate of $1\times 10^{-5}$, while the other randomly selects within $[5\times 10^{-6},1.5\times 10^{-5}]$ for each round. The conclusive results are depicted in Figure 4. Based on the experimental results, we observe that our method achieves faster loss convergence with the second and third types of signal quantities compared to the default and random settings, with the third type yielding the most impressive performance. In contrast, the first type of signal quantity had a negligible impact. These results suggest that while round-wise loss exhibits some degree of randomness, aggregating losses over multiple rounds can approximate heterogeneity to a meaningful extent, thus serving as an indicator to expedite model convergence.
It is also noteworthy that the third type of signal quantity aligns with the expression $\lVert\nabla f(\theta^{t})-\sum_{k=0}^{H}\eta_{i}\nabla f_{i}(\theta^{(t,k)})\rVert^{2}$, which most closely reflects Assumption 6. Consequently, it demonstrates the most effective performance in the experiments, not only achieving the fastest convergence but also the lowest stable loss. This case study experiment substantiates the efficacy of Proposition 3.7, offering valuable insights for parameter tuning in personalized FL.
6. Conclusion
This work investigates the convergence of FedMeZO, a practical approach integrating Memory-efficient Zeroth-Order Optimization within a federated learning setting for Large Language Models (LLMs). Extensive empirical results verified our analyses and indicated that FedMeZO achieves fast convergence with reduced GPU memory requirements, offering a promising alternative to traditional optimization methods. The incorporation of a personalized learning rate adjustment, derived from theoretical analysis, has been shown to effectively enhance loss reduction. Through this study, we aim to enlighten more research and development of memory-efficient optimization techniques to address practical challenges associated with the fine-tuning of LLMs, particularly in resource-constrained scenarios.
Appendix
Our appendix is organized as follows:
  • Section A summarizes all the used mathematical symbols.
  • Section B gives the introduced FL-related assumptions for our analysis.
  • Section C and Section D describe the detailed proof procedures of our lemmas and theorems respectively.
  • Section E details our empirical implementations in terms of the datasets, platforms, and hyper-parameters.
  • Section F presents additional experiment results about the conditions of learning rate (Section F.1), the impact of data heterogeneity (Section F.2), the impact of client number (Section F.3), the impact of perturbation scale (Section F.4), the impact of local iterations (Section F.5), and the full results of convergence curves (Section F.6).
Appendix A Notation
For ease of reading and reference, we present all mathematical symbols used in this paper in Table 2.
Table 2. Description of symbols used in the paper.
Symbol Description
$f(\theta)$ Global loss function over the parameter $\theta$
$\nabla f(\theta)$ Gradient of the loss function with respect to parameter $\theta$
$F_{i}(\theta,\mathcal{B}_{i})$ Local loss function on the $i^{th}$ client with mini-batch $\mathcal{B}_{i}$
$\nabla F(\theta,\mathcal{B})$ The gradient of parameter $\theta$ with mini-batches
~$\widetilde{\nabla}F_{i}(\theta,z,\mathcal{B}_{i},\mu)$ Zeroth-order gradient estimator for $F_{i}$ with perturbation $\mu$
$z$ Gaussian random variable sampled from $\mathcal{N}(0,I_{d})$
$\mathcal{B}_{i}$ Mini-batch of data sampled from the local distribution $\mathcal{D}_{i}$
$\mu$ Perturbation scale for zeroth-order gradient estimation
$\eta$ Learning rate for model updates
$n$ Number of perturbations in $n$-SPSA zeroth-order optimization
$(t,k)$ $k^{th}$ iteration within the $t^{th}$ communication round
$\mathcal{H}$ Hessian matrix of the loss function
$r$ The low effective rank of the Hessian matrix
$\gamma$ Factor quantifying the effective low-rank property of the gradient
$\zeta$ Factor quantifying the effective low-rank property of the gradient estimator
$d$ Dimension of the model parameter $\theta$
$L$ Smoothness constant of the loss function
$N$ Number of clients participating in FL
$i$ Index identifying the $i^{th}$ client
$T$ Number of communication rounds in FL
$e_{i}^{(t,k)}$ Zeroth-order gradient estimator of the $i^{th}$ client in iteration $(t,k)$
$H$ Total number of local iterations within a communication round
$c_{g}$, $\sigma_{g}$ Constants related to the gradient estimation gap caused by mini-batch stochasticity
$c_{h}$, $\sigma_{h}$ Constants related to the heterogeneity of client data and the global model
$\mathbb{E}_{t}$ Expectation taken over the randomness in the $t^{th}$ round
$\mathbb{E}_{t}^{k}$ Expectation taken over the randomness in the $k^{th}$ iteration of the $t^{th}$ round
Appendix B Detailed Assumptions
Assumption 2.
(Bounded Loss) The global loss function $f(\theta)$ is bounded below by a scalar $f^{*}$, i.e., $f^{*}\geq f(\theta)>-\infty$ for all $\theta$.
Assumption 3.
($L$-smoothness) The local and global loss functions $F_{i}(\theta,\mathcal{B}_{i})$, $f_{i}(\theta)$, and $f(\theta)$ are L-smooth. Mathematically, for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, it holds that
$\lVert\nabla f_{i}(\theta_{2})-\nabla f_{i}(\theta_{1})\rVert$ $\leq L\lVert\theta_{2}-\theta_{1}\rVert,$
$f_{i}(\theta_{2})\leq f_{i}(\theta_{1})+\langle\nabla f_{i}(\theta),\theta_{2}$ $-\theta_{1}\rangle+\frac{L}{2}{\lVert\theta_{2}-\theta_{1}\rVert}^{2}.$
Assumption 4.
(Mini-batch Gradient Error Bound) For any $\theta\in\mathbb{R}^{d}$, the second-order moment of the stochastic gradient is bounded by $\mathbb{E}_{\mathcal{B}_{i}}\lVert\nabla F_{i}(\theta,\mathcal{B}_{i})\rVert^{2}\leq c_{g}\lVert\nabla f_{i}(\theta)\rVert^{2}+\sigma_{g}^{2}$, where $c_{g}\geq 1$.
Assumption 5.
(Global-Local Disparities in i.i.d. Setting) For any $\theta\in\mathbb{R}^{d}$, the discrepancy between the local and global gradient is negligible, i.e., $\mathbb{E}_{i}\lVert\nabla f(\theta)-\nabla f_{i}(\theta)\rVert^{2}=0$.
Assumption 6.
(Global-Local Disparities in Non-i.i.d. Setting) For any $\theta\in\mathbb{R}^{d}$, the discrepancy between the local and global gradient is bounded by $\lVert\nabla f(\theta)-\nabla f_{i}(\theta)\rVert^{2}\leq c_{h}\lVert\nabla f(\theta)\rVert^{2}+\sigma_{h}^{2},$ where $c_{h}$ is a positive constant.
Assumptions 2-4 are well-established in the literature on large-scale stochastic optimization (Bottou et al., 2018). Assumption 5 describes an ideal i.i.d. setting where each client’s gradient is aligned with the global gradient. Assumption 6 accounts for the heterogeneity of client data distributions that is typical in non-i.i.d. settings (Li et al., 2019; Wang et al., 2021b; Li et al., 2020).
Appendix C Full Proof of Lemma
C.1. Proof of Lemma 2.2
We recap the proof of Stein’s identity following (Feng et al., 2023), noting that $\delta$ is a random variable sampled from $\mathcal{N}(0,\sigma^{2}I_{d})$, and we replace it with $\mu z_{i}$ in Definition 2.1.
$\nabla F_{i}(\theta,\mathcal{B}_{i})$ $=\nabla\mathbb{E}_{\delta\sim\mathcal{N}(0,\sigma^{2}I_{d})}\Bigl{[}F_{i}(\theta+\delta,\mathcal{B}_{i})\Bigr{]}$
[…]
$=\mathbb{E}_{\delta\sim\mathcal{N}(0,\sigma^{2}I_{d})}\Bigl{[}\frac{\delta}{\sigma^{2}}\cdot F_{i}(\theta+\delta,\mathcal{B}_{i})\Bigr{]}$
(28) $=\mathbb{E}_{\mu z_{i}\sim\mathcal{N}(0,\mu^{2}I_{d})}\Bigl{[}\frac{z_{i}}{\mu}\cdot F_{i}(\theta+\mu z_{i},\mathcal{B}_{i})\Bigr{]}.$
By symmetry, we change $\delta$ to $-\delta$ and obtain
(29) $\nabla F_{i}(\theta,\mathcal{B}_{i})$ $=-\mathbb{E}_{\mu z_{i}\sim\mathcal{N}(0,\mu^{2}I_{d})}\Bigl{[}\frac{z_{i}}{\mu}\cdot F_{i}(\theta-\mu z_{i},\mathcal{B}_{i})\Bigr{]}.$
Further, we prove that with
$\nabla F_{i}(\theta,\mathcal{B}_{i})$ $=\frac{1}{2}\mathbb{E}\Bigl{[}\frac{z_{i}}{\mu}\cdot F_{i}(\theta+\mu z_{i},\mathcal{B}_{i})\Bigr{]}-\frac{1}{2}\mathbb{E}\Bigl{[}\frac{z_{i}}{\mu}\cdot F_{i}(\theta-\mu z_{i},\mathcal{B}_{i})\Bigr{]}$
(30) $=\mathbb{E}\Bigl{[}\frac{z_{i}}{2\mu}\widetilde{\nabla}F_{i}(\theta,z_{i},\mathcal{B}_{i},\mu)\Bigr{]}.$
Finally, we adopt Eq. (1) and get Lemma 2.2, i.e.,
(31) $\nabla f_{i}(\theta)=\mathbb{E}[\nabla F_{i}(\theta,\mathcal{B}_{i})]=\mathbb{E}[\widetilde{\nabla}F_{i}(\theta,z_{i},\mathcal{B}_{i},\mu)].$
C.2. Proof of Lemma 2.3
Malladi et al. (Malladi et al., 2023) present a step-wise learning rate decay corollary that is independent of the dimensionality parameter $d$ and is solely related to the low effective rank $r$. We formalize this result as follows:
Lemma C.1 (Step-Wise Learning Rate Decay).
Assuming the Hessian matrix in terms of $\theta$ exhibits a local effective rank of $r$, and there exists a constant $\gamma=\frac{dr+d-2}{n(d+2)}=\Theta(r/n)$, the expected decrease in the loss can be bounded as follows:
(32) $\mathbb{E}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})-\frac{1}{\gamma}\eta\lVert\nabla f(\theta^{t})\rVert^{2}+\frac{1}{2}\eta^{2}L\frac{1}{\gamma}\mathbb{E}\bigl{[}{\lVert\nabla{F(\theta,\mathcal{B})}\rVert}^{2}\bigr{]},$
where $n$ denotes the number of randomizations.
It is important to note that the last term in Eq. (32) represents the squared norm of the true gradient, which is inconsistent with the zeroth-order estimation method used in our approach. Therefore, a transformation is necessary to align it with the FedMeZO algorithm. Besides, Malladi et al. detail the relationship between the true gradient norm and the zeroth-order gradient estimator norm, which we restate as follows:
Lemma C.2 (Gradient Estimator Norm Relationship).
The squared norm of the gradient estimated by the MeZO is given by
(33) $\mathbb{E}\Bigl{[}{\lVert\nabla{F(\theta,\mathcal{B})}\rVert}^{2}\Bigr{]}=\frac{n}{d+n-1}\mathbb{E}\Bigl{[}{\lVert\widetilde{\nabla}{F(\theta,z,\mathcal{B},\mu)}\rVert}^{2}\Bigr{]}.$
By substituting Eq. (33) into Eq. (32), we obtain Lemma 2.3. ∎
Appendix D Full Proof of Theorems
D.1. Proof of Theorem 3.1
For the term $T_{1}$:
$T_{1}$ $=\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}(\theta^{t})\Bigg{\rVert}}^{2}$
$=\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})+\nabla f(\theta^{t})\Bigr{]}\Bigg{\rVert}}^{2}$
$\leq 2\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})\Bigr{]}\Bigg{\rVert}}^{2}+2\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\nabla f(\theta^{t})\Bigg{\rVert}}^{2}$
$\leq\frac{2}{N^{2}}\sum_{i=1}^{N}\mathbb{E}_{t}{\Big{\lVert}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})\Big{\rVert}}^{2}+2\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}$
(34) $=2\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2},$
where the first inequality follows from the Cauchy-Schwarz inequality, the second inequality follows from Jensen’s inequality, and the third equality is derived from Assumption 5.
For the term $T_{2}$, by applying Jensen’s inequality, we obtain:
(35) $T_{2}\leq\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}{\Big{\lVert}e_{i}^{(t,k)}\Big{\rVert}}^{2},$
where $e_{i}^{(t,k)}$ represents the gradient estimator defined in Eq. (2.3). Substituting the estimator into the above expression and simplifying it, we obtain the following inequality. For brevity, let $\theta_{i}^{{}^{\prime}}$ denote $\theta_{i}^{(t,k)}$, $z_{i}^{{}^{\prime}}$ denote $z_{i}^{(t,k)}$, and $\mathcal{B}i^{{}^{\prime}}$ denote $\mathcal{B}i^{(t,k)}$:
$T_{2}$ $\leq\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}{\Bigg{\lVert}\frac{z_{i}^{{}^{\prime}}}{2\mu}(F_{i}(\theta_{i}^{{}^{\prime}}+\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}}-\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigg{\rVert}}^{2}$
$=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Bigg{\lVert}\frac{z_{i}^{{}^{\prime}}}{2\mu}\Bigl{(}F_{i}(\theta_{i}^{{}^{\prime}}+\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigr{)}$
$\hskip 99.58464pt+\Bigl{(}F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}}-\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigr{)}\Bigg{\rVert}^{2}$
$\leq\frac{2}{N^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Bigg{\lVert}\frac{z_{i}^{{}^{\prime}}}{2\mu}\Bigl{(}F_{i}(\theta_{i}^{{}^{\prime}}+\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigr{)}\Bigg{\rVert}^{2}$
(36) $\hskip 14.22636pt+\frac{2}{N^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Bigg{\lVert}\frac{z_{i}^{{}^{\prime}}}{2\mu}\Bigl{(}F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}}-\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigr{)}\Bigg{\rVert}^{2},$
where the inequality follows from the fact that $(\lVert a+b\rVert)^{2}\leq 2(\lVert a\rVert)^{2}+2(\lVert b\rVert)^{2}$. Due to the symmetry of the function $F_{i}$ when perturbed with Gaussian-distributed $z_{i}^{{}^{\prime}}$, both terms on the right-hand side are equivalent. Hence $T_{2}$ can be transformed into
$T_{2}$ $\leq\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Bigg{\lVert}\frac{z_{i}^{{}^{\prime}}}{\mu}\Bigl{(}F_{i}(\theta_{i}^{(t,k)}+\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigr{)}\Bigg{\rVert}^{2}$
$=\frac{1}{N^{2}\cdot d^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Bigg{\lVert}\frac{d\cdot z_{i}^{{}^{\prime}}}{\mu}\Bigl{(}F_{i}(\theta_{i}^{{}^{\prime}}+\mu z_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})-F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Bigr{)}\Bigg{\rVert}^{2}$
$\leq\frac{1}{N^{2}\cdot d^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\Biggl{[}2d\cdot\mathbb{E}_{t}\Big{\lVert}\nabla F_{i}(\theta_{i}^{{}^{\prime}},\mathcal{B}_{i}^{{}^{\prime}})\Big{\rVert}^{2}+\frac{\mu^{2}}{2}L^{2}d^{2}\Biggr{]}$
(37) $\leq\frac{1}{N^{2}\cdot d^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\Biggl{[}2c_{g}d\cdot\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{{}^{\prime}})\Big{\rVert}^{2}+2d\sigma_{g}^{2}+\frac{\mu^{2}}{2}L^{2}d^{2}\Biggr{]},$
where the first inequality follows Lemma 4.1 from (Gao et al., 2018), and the second inequality follows the Assumption 4. For the expectation term, we have
$\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}$ $=\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\mp\nabla f_{i}(\theta^{t})\mp\nabla f(\theta_{i}^{t})\Big{\rVert}^{2}$
(38) $\leq 3L^{2}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+3\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2},$
where the inequality is due to the Cauchy-Schwartz inequality, $L$-smooth property and Assumption 5. By taking Eq. (D.1) into Eq. (D.1), finally we can bound $T_{2}$ as:
$T_{2}\leq\frac{6c_{g}L^{2}}{N^{2}d}$ $\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$
(39) $+\frac{6c_{g}H}{Nd}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+\frac{2H\sigma_{g}^{2}}{Nd}+\frac{\mu^{2}HL^{2}}{2N}.$
After combining Eq. (4), Eq. (D.1) and Eq. (D.1), we have
$\mathbb{E}_{t}$ $\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})-\frac{2}{\gamma}\eta\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+\frac{3c_{g}\zeta\eta^{2}HL}{Nd}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$+\frac{3c_{g}\zeta\eta^{2}L^{3}}{N^{2}d}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}$
$=f(\theta^{t})+\Biggl{(}\frac{3c_{g}\zeta\eta^{2}HL}{Nd}-\frac{2}{\gamma}\cdot\eta\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
(40) $+\frac{3c_{g}\zeta\eta^{2}L^{3}}{N^{2}d}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}.$
Next we need to bound $\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$ and simplify Eq. (D.1). Specifically, by denoting $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{t}^{k-1}\lVert\theta_{i}^{(t,k)}-\theta^{t}\rVert^{2}$ as $s^{(t,k)}$, we have
$s^{(t,\tau)}$ $=\eta^{2}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{t}^{\tau-1}\Bigg{\lVert}\sum_{k=1}^{\tau}e_{i}^{(t,k)}\Bigg{\rVert}^{2}$
(41) $\leq\tau\eta^{2}\sum_{k=1}^{\tau}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{t}^{k}\Big{\lVert}e_{i}^{(t,k)}\Big{\rVert}^{2}.$
By combing Eq. (D.1), Eq. (D.1) and Eq. (D.1), we have
$s^{(t,\tau)}\leq 6c_{g}dL^{2}\tau\eta^{2}\sum_{k=1}^{\tau}s^{(t,k)}+6$ $c_{g}d\tau^{2}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$
(42) $+2d\sigma_{g}^{2}\tau^{2}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}\tau^{2}\eta^{2}}{2}.$
By taking summation over $\tau$ from 2 to $H$, and utilizing the property of arithmetic sequence, we obtain
$\sum_{\tau=2}^{H}s^{(t,\tau)}$ $\leq 6c_{g}dL^{2}\eta^{2}\sum_{\tau=2}^{H}\tau\sum_{k=1}^{\tau}s^{(t,k)}+C_{1}$
(43) $\leq 3c_{g}dH^{2}L^{2}\eta^{2}\sum_{k=1}^{H}s^{(t,k)}+C_{1},$
[…]
As $s^{(t,1)}=0$, after rearranging Eq. (D.1), we have
(44) $(1-3c_{g}dH^{2}L^{2}\eta^{2})\sum_{k=1}^{H}s^{(t,k)}\leq C_{1}.$
For simplification, here we denote $(1-3c_{g}dH^{2}L^{2}\eta^{2})$ as $C_{0}$. When $\eta\leq\frac{1}{3HL\sqrt{c_{g}d}}$, $C_{0}\geq\frac{2}{3}$. Under this condition, we have:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$ $=\sum_{k=1}^{H}s^{(t,k)}\leq\frac{C_{1}}{C_{0}}$
$=\frac{1}{C_{0}}\Biggl{(}2c_{g}dH^{3}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$ $+\frac{2}{3}d\sigma_{g}^{2}H^{3}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}H^{3}\eta^{2}}{6}\Biggr{)}$
(45) $\leq 3c_{g}dH^{3}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$ $+d\sigma_{g}^{2}H^{3}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}H^{3}\eta^{2}}{4}.$
Taking Eq. (D.1) into Eq. (D.1), we obtain the final result of loss descent of each step:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})+\Biggl{(}\frac{3c_{g}\zeta\eta^{2}HL}{Nd}-\frac{2}{\gamma}\cdot\eta\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\quad+\frac{3c_{g}\zeta\eta^{2}L^{3}}{Nd}\cdot\frac{C_{1}}{C_{0}}+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}$
$\leq f(\theta^{t})+\Biggl{(}\frac{3c_{g}\zeta\eta^{2}HL}{Nd}-\frac{2}{\gamma}\cdot\eta+\frac{9c_{g}^{2}\zeta\eta^{4}H^{3}L^{3}}{N}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\quad+\frac{3c_{g}\sigma_{g}^{2}\zeta\eta^{4}H^{3}L^{3}}{N}+\frac{3c_{g}d\zeta\eta^{4}\mu^{2}H^{3}L^{5}}{4N}+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}$
$\leq f(\theta^{t})+\Biggl{(}\frac{3c_{g}\zeta\eta^{2}HL}{Nd}+3c_{g}d\eta^{2}H^{2}L^{2}\frac{3c_{g}\zeta\eta^{2}HL}{Nd}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\quad-\frac{2\eta}{\gamma}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+3c_{g}d\eta^{2}H^{2}L^{2}\cdot\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}$
$\quad+3c_{g}d\eta^{2}H^{2}L^{2}\cdot\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}$
$\leq f(\theta^{t})+\Biggl{(}(1+3c_{g}dH^{2}L^{2}\eta^{2})\frac{3c_{g}\zeta\eta^{2}HL}{Nd}-\frac{2}{\gamma}\cdot\eta\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
(46) $\quad+(1+3c_{g}dH^{2}L^{2}\eta^{2})\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+(1+3c_{g}dH^{2}L^{2}\eta^{2})\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}.$
With the condition $\eta\leq\frac{1}{3HL\sqrt{c_{g}d}}$, we have $(1+3c_{g}dH^{2}L^{2}\eta^{2})\leq 2$. Then taking the condition $\eta\leq\frac{N}{3c_{g}HL}$ and $\eta\leq\frac{1}{H^{2}}$, we can transform Eq. (D.1) into the result of Theorem 3.1 as:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})+2\Biggl{(}\frac{\zeta}{d}\eta$ $-\frac{1}{\gamma}\eta\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
(47) $\qquad+\frac{2\sigma_{g}^{2}\zeta\eta L}{NHd}+\frac{\zeta\eta\mu^{2}L^{3}}{2NH}.$
D.2. Proof of Corollary 3.2
To get the result of global convergence, we rearrange Eq. (D.1) and similarly bound $\eta$ as $\eta\leq\frac{1}{H^{2}}$. For simplicity, we denotes $\frac{d-\zeta\gamma}{d\gamma}$ as $\Gamma$ and have
$2\eta\Gamma\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq f(\theta^{t})-\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}+\frac{2\sigma_{g}^{2}\zeta L}{NHd}\eta+\frac{\zeta\mu^{2}L^{3}}{2NH}\eta$
(48) $\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq\frac{f(\theta^{t})-\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}}{2\eta\Gamma}+\frac{\sigma_{g}^{2}\zeta L}{\Gamma NHd}+\frac{\zeta\mu^{2}L^{3}}{4\Gamma NH}.$
Notice that $\gamma=\Theta(r)$ and $\zeta=\Theta(\frac{1}{rd})$, since parameter $d$ is a large number, $d-\zeta\gamma=d-\Theta(\frac{1}{d})=d$, hence the dominant term of $\Gamma$ is $\frac{1}{\gamma}$, which follows $\Gamma=\Theta(\frac{1}{r})$. Then simultaneously summing over $T$ rounds on both sides and taking the average:
(49) $\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}\leq\frac{f(\theta^{0})-\mathbb{E}_{t}\bigl{[}f(\theta^{T})\bigr{]}}{2\eta T\Gamma}+\frac{\sigma_{g}^{2}\zeta L}{\Gamma NHd}+\frac{\zeta\mu^{2}L^{3}}{4\Gamma NH}.$
D.3. Proof of Theorem 3.4
Starting from Eq. (4), under the assumption of global-local dissimilarity (Assumption 6), $T_{1}$ becomes:
$T_{1}$ $=\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}(\theta^{t})\Bigg{\rVert}}^{2}$
$=\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})+\nabla f(\theta^{t})\Bigr{]}\Bigg{\rVert}}^{2}$
$\leq 2\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})\Bigr{]}\Bigg{\rVert}}^{2}+2\mathbb{E}_{t}{\Bigg{\lVert}\frac{1}{N}\sum_{i=1}^{N}\nabla f(\theta^{t})\Bigg{\rVert}}^{2}$
$\leq\frac{2}{N^{2}}\sum_{i=1}^{N}\mathbb{E}_{t}{\Big{\lVert}\nabla f_{i}(\theta^{t})-\nabla f(\theta^{t})\Big{\rVert}}^{2}+2\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}$
$\leq\frac{2}{N^{2}}\sum_{i=1}^{N}\mathbb{E}_{t}\Bigl{[}c_{h}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}+\sigma_{h}^{2}\Bigr{]}+2\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}$
(50) $=\frac{2(N+c_{h})}{N}\mathbb{E}_{t}{\big{\lVert}\nabla f(\theta^{t})\big{\rVert}}^{2}+\frac{2\sigma_{h}^{2}}{N},$
where the first inequality follows Cauchy-Schwartz, the second inequality follows Jensen’s inequality, and the third inequality follows Assumption 6.
Then we begin to bound the term $T_{2}$. Taking the result from Eq. (D.1), we have:
(51) $T_{2}\leq\frac{1}{N^{2}\cdot d^{2}}\sum_{i=1}^{N}\sum_{k=1}^{H}\Biggl{[}2c_{g}d\cdot\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}+2d\sigma_{g}^{2}+\frac{\mu^{2}}{2}L^{2}d^{2}\Biggr{]},$
where the first inequality follows Lemma 4.1 from (Gao et al., 2018), and the second inequality follows the Assumption 4. The term $\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}$ in Eq. (51) can be bounded as:
$\mathbb{E}_{t}\Big{\lVert}\nabla$ $f_{i}(\theta_{i}^{(t,k)})\Big{\rVert}^{2}=\mathbb{E}_{t}\Big{\lVert}\nabla f_{i}(\theta_{i}^{(t,k)})\mp\nabla f_{i}(\theta^{t})\mp\nabla f(\theta_{i}^{t})\Big{\rVert}^{2}$
(52) $\leq 3L^{2}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+3(c_{h}+1)\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+3\sigma_{h}^{2},$
where the inequality follows the Cauchy-Schwartz, $L$-smooth and Assumption 6. By applying Eq. (D.3) into Eq. (51), we obtain:
$T_{2}$ $\leq\frac{6c_{g}L^{2}}{N^{2}d}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+\frac{6c_{g}(c_{h}+N)H}{Nd}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
(53) $\hskip 85.35826pt+\frac{6c_{g}\sigma_{h}^{2}H}{Nd}+\frac{2H\sigma_{g}^{2}}{Nd}+\frac{\mu^{2}HL^{2}}{2N}.$
Combining Eq. (4), Eq. (D.3) and Eq. (D.3), we have:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})-\frac{2}{\gamma N}(N+c_{h})\eta\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}-\frac{2}{\gamma N}\sigma_{h}^{2}\eta$
$\quad+\frac{3c_{g}(c_{h}+N)\zeta\eta^{2}HL}{Nd}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+\frac{3c_{g}\sigma_{h}^{2}\zeta\eta^{2}HL}{Nd}$
$\quad+\frac{3c_{g}\zeta\eta^{2}L^{3}}{N^{2}d}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}$
$=f(\theta^{t})+\Biggl{(}\frac{3c_{g}(c_{h}+N)\zeta\eta^{2}HL}{Nd}-\frac{2}{\gamma N}\cdot(N+c_{h})\eta\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\quad+\frac{3c_{g}\zeta\eta^{2}L^{3}}{N^{2}d}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}+\frac{3c_{g}\sigma_{h}^{2}\zeta\eta^{2}HL}{Nd}$
(54) $\quad+\frac{\sigma_{g}^{2}\zeta\eta^{2}HL}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}-\frac{2}{\gamma N}\cdot\sigma_{h}^{2}\eta.$
Similar to Section D.1, now we bound $\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$ as follows:
$s^{(t,\tau)}$ $=\eta^{2}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{t}^{\tau-1}\Bigg{\lVert}\sum_{k=1}^{\tau}e_{i}^{(t,k)}\Bigg{\rVert}^{2}$
(55) $\leq\tau\eta^{2}\sum_{k=1}^{\tau}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{t}^{k}\Big{\lVert}e_{i}^{(t,k)}\Big{\rVert}^{2},$
where $s^{(t,k)}$ denotes $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{t}^{k-1}\lVert\theta_{i}^{(t,k)}-\theta^{t}\rVert^{2}$.
By combing Eq. (51), Eq. (D.3) and Eq. (D.3), we have
$s^{(t,\tau)}$ $\leq 6c_{g}dL^{2}\tau\eta^{2}\sum_{k=1}^{\tau}s^{(t,k)}+6c_{g}d(c_{h}+1)\tau^{2}\eta^{2}\mathbb{E}_{t}\big{\lVert}\nabla f(\theta^{t})\big{\rVert}^{2}$
(56) $\hskip 42.67912pt+6c_{g}d\sigma_{h}^{2}\tau^{2}\eta^{2}+2d\sigma_{g}^{2}\tau^{2}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}\tau^{2}\eta^{2}}{2}.$
By taking summation over $\tau$ from 2 to $H$, and utilizing the property of arithmetic sequence, we obtain
$\sum_{\tau=2}^{H}s^{(t,\tau)}$ $\leq 6c_{g}dL^{2}\eta^{2}\sum_{\tau=2}^{H}\tau\sum_{k=1}^{\tau}s^{(t,k)}+C_{2}$
(57) $\leq 3c_{g}dH^{2}L^{2}\eta^{2}\sum_{k=1}^{H}s^{(t,k)}+C_{2}$
[…]
$+2c_{g}$ $d\sigma_{h}^{2}H^{3}\eta^{2}+\frac{2}{3}d\sigma_{g}^{2}H^{3}\eta^{2}+\frac{\mu^{2}L^{2}d^{2}H^{3}\eta^{2}}{6}.$
Rearranging Eq. (D.3) and using $s^{(t,1)}=0$, we have
(58) $(1-3c_{g}dH^{2}L^{2}\eta^{2})\sum_{k=1}^{H}s^{(t,k)}\leq C_{2}.$
Let $C_{0}$ be $(1-3c_{g}dH^{2}L^{2}\eta^{2})$. When $\eta\leq\frac{1}{3HL\sqrt{c_{g}d}}$, $C_{0}\geq\frac{2}{3}$. Under this condition, we have:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{H}\mathbb{E}_{t}\Big{\lVert}\theta_{i}^{(t,k)}-\theta^{t}\Big{\rVert}^{2}$ $=\sum_{k=1}^{H}s^{(t,k)}\leq\frac{C_{2}}{C_{0}}$
$\leq 3c_{g}(c_{h}+1)d\eta^{2}H^{3}\mathbb{E}_{t}\big{\lVert}$ $\nabla f(\theta^{t})\big{\rVert}^{2}+3c_{g}\sigma_{h}^{2}d\eta^{2}H^{3}$
(59) $+\sigma_{g}^{2}d\eta^{2}H^{3}+\frac{d^{2}\eta^{2}\mu^{2}H^{3}L^{2}}{4}.$
Let $\widetilde{c}_{h}$ be $(c_{h}+N)$ and $\widetilde{\sigma}^{2}$ be $(3c_{g}\sigma_{h}^{2}+\sigma_{g}^{2})$. Applying Eq. (D.3) into Eq. (D.3), we can obtain the final result of stepwise loss descent:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}\leq f(\theta^{t})+\frac{3\zeta\eta^{2}LHc_{g}(c_{h}+N)}{Nd}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\qquad-\frac{2}{\gamma N}\cdot\eta(N+c_{h})\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+\frac{3\zeta\eta^{2}L^{3}c_{g}}{Nd}\cdot\frac{C_{2}}{C_{0}}$
$\qquad+\frac{3\zeta\eta^{2}c_{g}\sigma_{h}^{2}HL}{Nd}+\frac{\zeta\eta^{2}H\sigma_{g}^{2}L}{Nd}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}-\frac{2}{\gamma N}\cdot\eta\sigma_{h}^{2}$
$\leq f(\theta^{t})+\Biggl{(}\frac{3\zeta\eta^{2}LHc_{g}\widetilde{c}_{h}}{Nd}-\frac{2\eta\widetilde{c}_{h}}{\gamma N}+\frac{9\zeta c_{g}^{2}\widetilde{c}_{h}H^{3}L^{3}\eta^{4}}{N}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\qquad+\frac{3c_{g}\widetilde{\sigma}^{2}\zeta H^{3}L^{3}\eta^{4}}{N}+\frac{3c_{g}\zeta d\mu^{2}H^{3}L^{5}\eta^{4}}{4N}+\frac{\widetilde{\sigma}^{2}\zeta HL\eta^{2}}{Nd}$
$\qquad+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}-\frac{2}{\gamma N}\cdot\eta\sigma_{h}^{2}$
$\leq f(\theta^{t})+\Biggl{(}(1+3c_{g}dH^{2}L^{2}\eta^{2})\frac{3\zeta\eta^{2}LHc_{g}\widetilde{c}_{h}}{Nd}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\qquad-\frac{2}{\gamma N}\eta\widetilde{c}_{h}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}+3c_{g}dH^{2}L^{2}\eta^{2}\frac{\widetilde{\sigma}^{2}}{Nd}\zeta HL\eta^{2}-\frac{2}{\gamma N}\eta\sigma_{h}^{2}$
$\qquad+3c_{g}dH^{2}L^{2}\eta^{2}\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}+\frac{\widetilde{\sigma}^{2}}{Nd}\zeta HL\eta^{2}+\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}$
$\leq f(\theta^{t})+\Biggl{(}(1+3c_{g}dH^{2}L^{2}\eta^{2})\frac{3\zeta\eta^{2}LHc_{g}\widetilde{c}_{h}}{Nd}-\frac{2\eta\widetilde{c}_{h}}{\gamma N}\Biggr{)}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
$\qquad+(1+3c_{g}dH^{2}L^{2}\eta^{2})\cdot\frac{\widetilde{\sigma}^{2}}{Nd}\zeta HL\eta^{2}$
(60) $\qquad+(1+3c_{g}dH^{2}L^{2}\eta^{2})\cdot\frac{\zeta\eta^{2}\mu^{2}HL^{3}}{4N}-\frac{2}{\gamma N}\cdot\eta\sigma_{h}^{2}.$
With the condition $\eta\leq\frac{1}{3HL\sqrt{c_{g}d}}$, we have $(1+3c_{g}dH^{2}L^{2}\eta^{2})\leq 2$. Taking $\eta\leq\frac{N}{3c_{g}HL}$, $\eta\leq\frac{1}{H^{2}}$, Eq. (D.3) becomes:
$\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}$ $\leq f(\theta^{t})+2\Biggl{(}\frac{\zeta}{d}-\frac{1}{\gamma N}\Biggr{)}\widetilde{c}_{h}\eta\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$
(61) $\qquad+\frac{2\widetilde{\sigma}^{2}\zeta L\eta}{NHd}+\frac{\zeta\eta\mu^{2}L^{3}}{2NH}-\frac{2}{\gamma N}\cdot\eta\sigma_{h}^{2}.$
D.4. Proof of Corollary 3.2
Denote $\widetilde{\Gamma}=\frac{d-N\gamma\zeta}{d\gamma N}$. Rearranging Eq. (D.3), simultaneously summing over $T$ rounds on both sides and taking the average, we get the result:
$2\widetilde{\Gamma}\widetilde{c}_{h}\eta\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq f(\theta^{t})-\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}$
$\qquad+\frac{2\widetilde{\sigma}^{2}\zeta L\eta}{NHd}+\frac{\zeta\eta\mu^{2}L^{3}}{2NH}-\frac{2}{\gamma N}\cdot\eta\sigma_{h}^{2}$
$\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq\frac{f(\theta^{t})-\mathbb{E}_{t}\bigl{[}f(\theta^{t+1})\bigr{]}}{2\widetilde{\Gamma}\widetilde{c}_{h}\eta}$
$\qquad+\frac{\widetilde{\sigma}^{2}\zeta L}{\widetilde{\Gamma}\widetilde{c}_{h}NHd}+\frac{\zeta\mu^{2}L^{3}}{\widetilde{\Gamma}\widetilde{c}_{h}4NH}-\frac{\sigma_{h}^{2}}{\widetilde{\Gamma}\widetilde{c}_{h}\gamma N}$
$\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}_{t}\Big{\lVert}\nabla f(\theta^{t})\Big{\rVert}^{2}$ $\leq\frac{f(\theta^{0})-f^{*}}{2\widetilde{\Gamma}\widetilde{c}_{h}\eta T}+\frac{\widetilde{\sigma}^{2}\zeta L}{\widetilde{\Gamma}\widetilde{c}_{h}NHd}$
(62) $\qquad+\frac{\zeta\mu^{2}L^{3}}{\widetilde{\Gamma}\widetilde{c}_{h}4NH}-\frac{\sigma_{h}^{2}}{\widetilde{\Gamma}\widetilde{c}_{h}\gamma N}.$
Appendix E Implementation Details
In this section, we provide the detailed implementations of our experiments. Some experimental settings have already been discussed in Section 5.1 and will not be reiterated here.
E.1. Datasets
Table 3. Datasets and Basic Information.
Name #Sample Domain
Fed-Alpaca 52.0k Generic Language
Fed-Dolly 15.0k Generic Language
Fed-GSM8K 7.5k CoT
Fed-CodeAlpaca 8.0k Code Generation
We adopt several federated tuning datasets tailored for LLMs from (Kuang et al., 2023), with different splitting strategies to simulate the heterogeneity typical of different federated learning (FL) scenarios, including a uniform distribution of data, a Dirichlet distribution of data, and a splitter based on meta-information.
Fed-Dolly: This federated corpus dataset, derived from Databricks-dolly-15k (Conover et al., 2023a), comprises eight categories of NLP tasks: brainstorming, classification, closed QA, creative writing, general QA, information extraction, open QA, and summarization. We divide the training set into three subsets using a three-way split and assign each subset to a distinct client.
Fed-GSM8K: Constructed from the GSM8K dataset (Cobbe et al., 2021), this collection is aimed at mathematical fine-tuning and consists of 7.5K training problems alongside 1K test problems. By default, we partition the training set uniformly into three subsets and allocate each to a separate client.
Fed-CodeAlpaca: This federated version of CodeAlpaca (Chaudhary, 2023) encompasses code samples in ten programming languages, including C, C#, C++, Go, Java, PHP, Pascal, Python, Scala, and X86-64 Assembly. Due to the scarcity of X86-64 Assembly samples in the original corpus, we exclude them. The remaining samples are then divided into three subsets using a default three-way split, with one subset assigned to each client.
Fed-Alpaca: The Alpaca dataset (Taori et al., 2023) is designed for LLM fine-tuning and features natural language questions and responses for a variety of NLP tasks, such as text generation, translation, and open QA. It spans various domains like math, text processing, and code generation.
E.2. Experimental Platforms
We implement our approaches using PyTorch (Paszke et al., 2019) v1.10.1, coupled with PEFT v0.3.0 and the Transformers library (Wolf et al., 2020) v4.29.2. Experiments with LLaMA-3B are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers.
E.3. Default Implementation Settings
Following the guidelines in (Kuang et al., 2023; Malladi et al., 2023), all approaches perform local training with a batch size of 1 to minimize memory usage. In an effort to standardize the experimental conditions, both backpropagation (BP)-based methods and our proposed method FedMeZO, train locally with specific learning rates: $\eta=1\times 10^{-5}$ for the Fed-Dolly and Fed-Alpaca datasets, $\eta=2\times 10^{-5}$ for the Fed-CodeAlpaca dataset, and $\eta=2.5\times 10^{-5}$ for the Fed-GSM8K dataset. The rank and alpha parameters for Low-Rank Adaptation (LoRA) adapters used by both BP-based optimization and FedMeZO are set to 128 and 256, respectively. As per (Malladi et al., 2023), the perturbation scale $\mu$ for FedMeZO is set to $1\times 10^{-3}$. Unless otherwise stated, in our training process, we employed the early stopping mechanism to prevent over-fitting and reduce unnecessary training time caused following previous works (Chen et al., 2022). The training was stopped if there was no improvement in the validation loss for a predefined number of consecutive epochs, known as the patience parameter. We chose a patience of 30 epochs based on empirical evidence or prior studies. The best model was selected from the epoch with the lowest validation loss. Furthermore, apart from individual experiments with time constraints (experiments in Appendix F.1 and Appendix F.3), we conducted three sets of experiments with randomly selected seeds for the same set of parameters, and calculated the mean as the line plot with a 90% confidence interval as the error bar.
The influence of different hyper-parameters for FedMeZO has been analyzed in Section 5.
Appendix F Supplementary Experiments
F.1. Conditions for the Learning Rate
To validate the theory regarding the recommended learning rate settings mentioned in Section 3.3, we conducted the following experiment. Firstly, we disabled the early-stop mechanism to observe the subsequent changes brought about by larger learning rates. Secondly, on the GSM-IID dataset, we sequentially selected four learning rates: $1\times 10^{-5}$, $3\times 10^{-5}$, $5\times 10^{-5}$, and $1\times 10^{-4}$, and conducted training for 500 rounds each. The final results are presented in Figure 5.
The results indicate that when the learning rate exceeds the range supported by theory, the loss function exhibits a sharp increase, and the larger the learning rate, the earlier this sharp increase occurs.
It is noteworthy that our theory suggests that an optimal learning rate magnitude is anchored at $1/\sqrt{d}$ in Section 3.3. In our chosen model, LLaMA-3B, where $d$ can be set to $3\times 10^{9}$, this gives $1/\sqrt{d}=1.826\times 10^{-5}$, which is approximately the learning rate we aim to use. Exceeding this learning rate might lead to unexpected outcomes.
Therefore, this finding underscores the importance of adhering to theoretical guidelines for setting learning rates in order to avoid destabilizing the training process and ensuring a smooth convergence toward the optimal solution.
Figure 5. Phenomenon of loss surge due to larger learning rates.
F.2. The Impact of Heterogeneity on Convergence
We present the results of processing the same dataset with different splitters in Figure 6. It is observable that in the Dolly and CodeAlpaca datasets, the LDA and Meta splitters perform better than IID, with Meta being the best. Noting that the data classified by Meta and LDA are non-i.i.d., this indicates that higher data heterogeneity is more conducive to model convergence.
Figure 6. Effects of different splitters on the same dataset.
F.3. The Impact of Client Number
In Figure 7, we showcase the outcomes of federated learning on the same Alpaca dataset with 3 clients and 8 clients, respectively. Initially, there is no significant difference between the two during the early rounds of training. However, after 200 rounds, the training loss with 3 clients has dropped to its lowest and begins to fluctuate, while the training loss with 8 clients continues to steadily converge. This demonstrates that the model converges more stably with more clients participating in the training. This conclusion also corresponds to the theoretical results regarding the number of clients $N$ discussed in Section 3, i.e., an increase in $N$ is beneficial for reducing global convergence.
Figure 7. Effects of the number of clients.
F.4. The Impact of Perturbation Scale
We display the comprehensive experimental results of altering the perturbation scale $\mu$ across different datasets in Figure 8. As mentioned in Section 5.3.1, we observe that, except for CodeAlpaca, smaller values of $\mu$ slightly accelerate the convergence speed in the remaining results, with their corresponding lines all positioned below the default setting, whereas larger values of $\mu$ result in slower convergence speeds. Through these extensive experiments, we further substantiate the theoretical findings.
Figure 8. Effects of different perturbation scales $\mu$.
F.5. The Impact of Local Iterations
We present the complete experimental results of changing the local iteration $H$ on different datasets in Figure 9. As mentioned in Section 5.3.2, we find that across all datasets and splitters, a larger $H$ significantly speeds up convergence, although in certain scenarios, such as Code-IID and Code-Meta, the loss does not reach a lower stable convergence state, and a smaller $H$ consistently results in slower convergence speeds. Through these comprehensive experiments, we further validate the theoretical results in Section 3 related to the impact of $H$.
Figure 9. Effects of different local iterations $H$.
F.6. Comprehensive Results of Convergence Study
As mentioned in Section 5.2, we have validated the convergence of BP-based FedAvg and FedMeZO across different datasets and splitters, with the results displayed in Figure 10. We observed that in all scenarios, FedMeZO achieves a faster loss reduction compared to BP-based FedAvg and attains a lower loss in certain datasets, such as Code-IID and GSM-IID. This indicates that under equivalent conditions, FedMeZO is more adept at learning the characteristics of different datasets. In the CodeAlpaca dataset, we observe that the BP-based FedAvg exhibits fluctuations with large error bars, showing a trend similar to that observed at the end of Code-Meta in Figure 9. We hypothesize that the increased fluctuation is attributable to the code data’s inherently disjointed nature in natural language terms, compared to datasets from other domains.
Figure 10. Convergence comparison of FedMeZO and BP-based FedAvg algorithm.